Is 3Fn-Fn-2=Fn+2for n >3n are valid?


1
Expert's answer
2022-04-13T18:11:35-0400

"3F_n-F_{n-2}=F_{n+2}\\F_{n+2}=F_{n+1}+F_n=F_{n-1}+F_n+F_n=\left( F_n-F_{n-2} \right) +F_n+F_n=3F_n-F_{n-2}\\True"